OpenStudy (anonymous):

5/7 x = 140 The x is? A.28 B.198 C.20 D.100

OpenStudy (anonymous):

can somebody help?

OpenStudy (gabbyalicorn):

57x=140 Multiply both sides by 7/5. (75)*(57x)=(75)*(140) x=196 Answer: x=196

OpenStudy (anonymous):

Thanks!

OpenStudy (gabbyalicorn):

No problem! ^-^